bond lenth of O22+ or nitrogen molecule is more

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Nidhi,

O22+

σ1s2<σ*1s2<σ2s2<σ*2s2<σ2px2*2px2<∏2py1=∏2pz1

bond order =2/2=1

N2

σ1s2<σ*1s2<σ2s2<σ*2s2<σ2px2<∏2py2=∏2pz2

bond order =6/2=3

therefore bond length of nitrogen molecule is less.